# Adaptive Deadline and Batch Layered Synchronized Federated Learning

### v1

@@ A.1 Appendix A - Proof of Lemma 1
$P(|\mathcal{U}_{t}^{l}=0|)=\left(P(z_{t}^{1}\leq L-l)\right)^{U}=\left(\sum_{k=0}^{L-l}\frac{\lambda_{t}^{k}e^{-\lambda_{t}}}{k!}\right)^{U}.$ (A.1)
For integer $s>1$, we perform integration by parts with the choices:
@@
[…]
@@
$Q(s,x)=\sum_{k=0}^{s-1}\frac{x^{k}e^{-x}}{k!}.$ (A.2)
Now, substituting Eq. (A.1) into Eq. (A.2), we get:
$p_{l}=P(|\mathcal{U}_{t}^{l}=0|)=Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U},$
proving Lemma 1. $\square$
A.2 Appendix B - Proof of Lemma 2
$\mathbb{E}\Big{[}\tilde{{\bm{w}}}_{t+1}^{l}\Big{]}=\mathbb{E}\Big{[}\mathbb{E}\Big{[}\tilde{{\bm{w}}}_{t+1}^{l}\Big{|}|\mathcal{U}_{t}^{l}|\Big{]}\Big{]}=\sum_{K=0}^{U}P(|\mathcal{U}_{t}^{l}|=K)\mathbb{E}\Big{[}\tilde{{\bm{w}}}_{t+1}^{l}\Big{|}|\mathcal{U}_{t}^{l}|=K\Big{]}.$
Substituting the probability of $|\mathcal{U}_{t}^{l}|=0$, denoted as $p_{l}$, we get:
$\mathbb{E}[\tilde{{\bm{w}}}_{t+1}^{l}]=p_{l}\tilde{{\bm{w}}}_{t}^{l}+\sum_{K=1}^{U}P(|\mathcal{U}_{t}^{l}|=K)\frac{1}{1-p_{l}}\left(\frac{1}{K}\mathbb{E}\left[\sum_{u\in\mathcal{U}_{t}^{l}}{\bm{w}}_{u,t}^{l}\middle||\mathcal{U}_{t}^{l}|=K\right]-p_{l}\tilde{{\bm{w}}}_{t}^{l}\right).$
$\mathbb{E}[\tilde{{\bm{w}}}_{t+1}^{l}]=p_{l}\tilde{{\bm{w}}}_{t}^{l}+\frac{{\bm{w}}_{t+1}^{l}-p_{l}\tilde{{\bm{w}}}_{t}^{l}}{1-p_{l}}\sum_{u=1}^{U}P(|\mathcal{U}_{t}^{l}|=K).$
Using:
$\sum_{u=1}^{U}P(|\mathcal{U}_{t}^{l}|=K)=1-P(|\mathcal{U}_{t}^{l}|=0)=1-p_{l},$
we get:
$\mathbb{E}[\tilde{{\bm{w}}}_{t+1}^{l}]=p_{l}\tilde{{\bm{w}}}_{t}^{l}+{\bm{w}}_{t+1}^{l}-p_{l}\tilde{{\bm{w}}}_{t}^{l}={\bm{w}}_{t+1}^{l}.$
Since this equality holds for every layer $l\in\{1,\dots,L\}$, and the global model $\tilde{{\bm{w}}}_{t+1}$ is formed by stacking the layer-wise vectors $\tilde{w}^{l}_{t+1}$, it follows that the full model update is also unbiased:
$\mathbb{E}[\tilde{{\bm{w}}}_{t+1}]={\bm{w}}_{t+1},$
completing the proof. $\square$
A.3 Appendix C - Proof of Lemma 3
$v_{t+1}^{l}\triangleq\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big{]}=P\Big{[}|\mathcal{U}_{t}^{l}|=0\Big{]}\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\ \Big{|}|{\mathcal{U}}_{t}^{l}|=0\Big{]}$ (A.4)
For the first summoned, it holds that:
$\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\ \Big{|}|{\mathcal{U}}_{t}^{l}|=0\Big{]}=\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big{]}=\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t}^{l}+{\bm{w}}_{t}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big{]}$ (A.5)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t}^{l}\|^{2}\Big{]}+\mathbb{E}\Big{[}\|{\bm{w}}_{t+1}^{l}-{\bm{w}}_{t}^{l}\|^{2}\Big{]}\stackrel{{\scriptstyle(b)}}{{\leq}}v_{t}^{l}+\sum_{u=1}^{U}\frac{1}{U}\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{u,t}^{l}-{\bm{w}}_{t}^{l}\|^{2}\Big{]}$
$=v_{t}^{l}+\sum_{u=1}^{U}\frac{1}{U}\|\eta_{t}\nabla F_{u}({\bm{w}}_{t};i_{t}^{u})\|^{2}\stackrel{{\scriptstyle(c)}}{{\leq}}v_{t}^{l}+\eta_{t}^{2}G^{2},$
where (a) is given by the triangle inequality, (b) from the convexity of $||\cdot||^{2}$ and (c) from Assumption A3.
For the second summoned:
$\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\ \Big{|}|{\mathcal{U}}_{t}^{l}|=K\Big{]}=\mathbb{E}\Bigg{[}\Bigg{|}\Bigg{|}\frac{1}{1-p_{l}}\left(\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-p_{l}\tilde{{\bm{w}}}_{t}^{l}\right)-{\bm{w}}_{t+1}^{l}\Bigg{|}\Bigg{|}^{2}\Bigg{|}|\mathcal{U}_{t}^{l}|=K\Bigg{]}$
$=\frac{1}{(1-p_{l})^{2}}\left(\mathbb{E}\Bigg{[}\Bigg{|}\Bigg{|}\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t+1}^{l}-p_{l}(\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l})\Bigg{|}\Bigg{|}^{2}\Bigg{|}|\mathcal{U}_{t}^{l}|=K\Bigg{]}\right)$
$=\frac{1}{(1-p_{l})^{2}}\times$
$\Bigg{(}\mathbb{E}\Bigg{[}\Bigg{|}\Bigg{|}\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t+1}^{l}\Bigg{|}\Bigg{|}^{2}\Bigg{|}|\mathcal{U}_{t}^{l}|=K\Bigg{]}$ (A.6)
$+p_{l}^{2}\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big{]}$ (A.7)
$-2p_{l}(\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l})^{R}\mathbb{E}\Bigg{[}\left(\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t+1}^{l}\right)\Bigg{|}|\mathcal{U}_{t}^{l}|=K\Bigg{]}\Bigg{)}.$ (A.8)
$\mathbb{E}\Bigg{[}\Bigg{|}\Bigg{|}\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t+1}^{l}\Bigg{|}\Bigg{|}^{2}\Bigg{|}|\mathcal{U}_{t}^{l}|=K\Bigg{]}\leq\frac{U}{K(U-1)}\left(1-\frac{K}{U}\right)4\eta_{t}^{2}G^{2}.$
$p_{l}^{2}\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big{]}\leq p_{l}^{2}\left(v_{t}^{l}+\eta_{t}^{2}G^{2}\right).$
$\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\ \Big{|}|{\mathcal{U}}_{t}^{l}|=K\Big{]}\leq\frac{1}{(1-p_{l})^{2}}\left(p_{l}^{2}(\eta_{t}^{2}G^{2}+v_{t}^{l})+\frac{U}{K(U-1)}\left(1-\frac{K}{U}\right)4\eta_{t}^{2}G^{2}\right).$ (A.9)
Plugging Eq. (A.5) and Eq. (A.9) into Eq. (A.4), we obtain:
$+\sum_{K=1}^{U}P\Big{[}|\mathcal{U}_{t}^{l}|=K\Big{]}\frac{1}{(1-p_{l})^{2}}\left(p_{l}^{2}(\eta_{t}^{2}G^{2}+v_{t}^{l})+\frac{U}{K(U-1)}\left(1-\frac{K}{U}\right)4\eta_{t}^{2}G^{2}\right)$
$\leq p_{l}\left(v_{t}^{l}+\eta_{t}^{2}G^{2}\right)+\frac{1}{(1-p_{l})^{2}}\left(p_{l}^{2}(\eta_{t}^{2}G^{2}+v_{t}^{l})+\frac{4\eta_{t}^{2}G^{2}U}{U-1}\right)\sum_{K=1}^{U}P\Big{[}|\mathcal{U}_{t}^{l}|=K\Big{]}$
$=p_{l}\left(v_{t}^{l}+\eta_{t}^{2}G^{2}\right)+\frac{1}{1-p_{l}}\left(p_{l}^{2}(\eta_{t}^{2}G^{2}+v_{t}^{l})+\frac{4\eta_{t}^{2}G^{2}U}{U-1}\right)$
$=\frac{p_{l}}{1-p_{l}}v_{t}^{l}+\eta_{t}^{2}G^{2}\left(\frac{\frac{4U}{U-1}+p_{l}}{1-p_{l}}\right).$
$v_{t+1}^{l}\leq D\cdot v_{t}^{l}+E,\quad v_{1}^{l}=0,$ (A.10)
where
$D\triangleq\frac{p_{l}}{1-p_{l}}\quad\text{and}\quad E\triangleq\eta_{t}^{2}G^{2}\left(\frac{\frac{4U}{U-1}+p_{l}}{1-p_{l}}\right).$ (A.11)
Since $p_{l}\leq p_{1}<0.5$, it follows that $D=\frac{p_{l}}{1-p_{l}}<1$, ensuring that the recurrence is contractive and thus leads to convergence of the variance sequence:
$v_{t+1}^{l}\leq D\cdot v_{t}^{l}+E\leq D^{2}v_{t-1}^{l}+E+DE\stackrel{{\scriptstyle(a)}}{{\leq}}E\sum_{k=0}^{t-1}D^{k}+\underbrace{D^{t}v_{1}^{l}}_{0}=\frac{1-D^{t}}{1-D}E,$
where (a) follows since Eq. (A.10).
Substituting $D$ and $E$ given in Eq. (A.11), we have:
$v_{t+1}^{l}\triangleq\sum_{l=1}^{L}\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big{]}\leq\frac{1-D^{t}}{1-D}E\leq\frac{E}{1-D}=\eta_{t}^{2}G^{2}\left(\frac{\frac{4U}{U-1}+p_{l}}{1-2p_{l}}\right).$
Now, summing over the $L$ layers:
$\leq\sum_{l=1}^{L}\eta_{t}^{2}G^{2}\left(\frac{\frac{4U}{U-1}+p_{l}}{1-2p_{l}}\right)\leq\eta_{t}^{2}G^{2}\frac{4U}{U-1}\sum_{l=1}^{L}\frac{1+p_{l}}{1-2p_{l}}.$
Substituting $p_{l}$, as given in Lemma 1, with the value in Lemma 1 gives:
$\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}\right\|^{2}\right]\leq\eta_{t}^{2}G^{2}\frac{4U}{(U-1)}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}{1-2Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}},$
thus proving Lemma 3. $\square$
A.4 Appendix D - Proof of Theorem 1
We start by using the full participation FedAvg update ${\bm{w}}_{t+1}$:
$\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{opt}\|^{2}\Big{]}$ $=\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}+{\bm{w}}_{t+1}-{\bm{w}}_{opt}\|^{2}\Big{]}$
$=\underbrace{\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}\|^{2}\Big{]}}_{S_{1}}+\underbrace{\mathbb{E}\Big{[}\|{\bm{w}}_{t+1}-{\bm{w}}_{opt}\|^{2}\Big{]}}_{S_{2}}+\underbrace{2\langle\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1},{\bm{w}}_{t+1}-{\bm{w}}_{opt}\rangle}_{S_{3}}.$
We note that $S_{3}$ vanishes by the unbiasedness Lemma 2. $S_{1}$ is bounded using Lemma 3:
$\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}\right\|^{2}\right]\leq\eta_{t}^{2}G^{2}\frac{4U}{(U-1)}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}{1-2Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}.$
The term $S_{2}$ is bounded using Lemma 1 from [6], which states that under Assumption A1 and a step size satisfying $\eta_{t}\leq\frac{1}{4L}$, the following inequality holds:
$\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{opt}\|^{2}\Big{]}=(1-\eta_{t}\rho_{c})\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{opt}\|^{2}\Big{]}+\underbrace{\eta_{t}^{2}\mathbb{E}\Big{[}\|g_{t}-\bar{g}_{t}\|^{2}\Big{]}}_{\tilde{F}}+6\eta_{t}^{2}\rho_{s}\Gamma,$ (A.12)
where $g_{t}$ and $\bar{g}_{t}$ are defined by:
$\eta_{t}^{2}\mathbb{E}\Big{[}\|g_{t}-\bar{g}_{t}\|^{2}\Big{]}$ $=\eta_{t}^{2}\mathbb{E}\Bigg{|}\Bigg{|}\sum_{u=1}^{U}\frac{1}{U}\left(F_{u}\left({\bm{w}};i_{t}^{u}\right)-\nabla F_{u}\left({\bm{w}}\right)\right)\Bigg{|}\Bigg{|}^{2}$
$=\eta_{t}^{2}\frac{1}{U^{2}}\sum_{u=1}^{U}\mathbb{E}\Big{|}\Big{|}\left(F_{u}\left({\bm{w}};i_{t}^{u}\right)-\nabla F_{u}\left({\bm{w}}\right)\right)\Big{|}\Big{|}^{2}$
$\leq\eta_{t}^{2}\frac{1}{mU^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{P_{u}},$
where the last inequality stems from Assumption A2.
Combining the bounds for $S_{1}$ and $S_{2}$, we obtain:
$\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{opt}\|^{2}\Big{]}\leq(1-\eta_{t}\rho_{c})\mathbb{E}\Big{[}\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{opt}\|^{2}\Big{]}+\eta_{t}^{2}\left(B+C_{t}\right),$
where $B$ and $C_{t}$ are given by:
$B\triangleq\frac{1}{mU^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{P_{u}}+6\rho_{s}\Gamma\quad C_{t}\triangleq G^{2}\frac{4U}{U-1}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}{1-2Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}.$
$\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{R+1}-{\bm{w}}_{opt}\right\|^{2}\right]\leq\prod_{t=1}^{R}\left(1-\eta_{t}\rho_{c}\right)\Delta_{1}+\sum_{t=1}^{R}\eta_{t}^{2}\left(B+C_{t}\right)\prod_{\tau=t+1}^{R}\left(1-\eta_{\tau}\rho_{c}\right),$

### v2

@@ A-C Appendix C - Proof of Lemma 3
$=$ $\mathbb{E}\Bigg[\Bigg|\Bigg|\frac{1}{1-p_{t}^{l}}\left(\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-p_{t}^{l}\tilde{{\bm{w}}}_{t}^{l}\right)-{\bm{w}}_{t+1}^{l}\Bigg|\Bigg|^{2}\Bigg||\mathcal{U}_{t}^{l}|=K\Bigg]$
$=$ $\frac{1}{(1-p_{t}^{l})^{2}}\times\Bigg(\mathbb{E}\Bigg[\Bigg|\Bigg|\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t+1}^{l}$
$-p_{t}^{l}(\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l})\Bigg|\Bigg|^{2}\Bigg||\mathcal{U}_{t}^{l}|=K\Bigg]\Bigg)$
$=$ $\frac{1}{(1-p_{t}^{l})^{2}}\times\Bigg(\mathbb{E}\Bigg[\Bigg|\Bigg|\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t+1}^{l}\Bigg|\Bigg|^{2}\Bigg||\mathcal{U}_{t}^{l}|=K\Bigg]$ (A.10)
$+(p_{t}^{l})^{2}\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big]-2p_{t}^{l}(\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l})^{T}$ (A.11)
$\cdot\mathbb{E}\Bigg[\left(\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t+1}^{l}\right)\Bigg||\mathcal{U}_{t}^{l}|=K\Bigg]\Bigg).$ (A.12)
$\mathbb{E}\Bigg[\Bigg|\Bigg|\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t+1}^{l}\Bigg|\Bigg|^{2}\Bigg||\mathcal{U}_{t}^{l}|=K\Bigg]$
$\leq\frac{U}{K(U-1)}\left(1-\frac{K}{U}\right)4\eta_{t}^{2}G^{2}.$
$(p_{t}^{l})^{2}\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big]\leq(p_{t}^{l})^{2}\left(v_{t}^{l}+\eta_{t}^{2}G^{2}\right).$
$\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\ \Big||{\mathcal{U}}_{t}^{l}|=K\Big]\leq\frac{1}{(1-p_{t}^{l})^{2}}\times$ (A.13)
$\left((p_{t}^{l})^{2}(\eta_{t}^{2}G^{2}+v_{t}^{l})+\frac{U}{K(U-1)}\left(1-\frac{K}{U}\right)4\eta_{t}^{2}G^{2}\right).$
Plugging Eq. (A.9) and Eq. (A.13) into Eq. (A.8), we obtain:
$+\sum_{K=1}^{U}P\Big[|\mathcal{U}_{t}^{l}|=K\Big]\frac{1}{(1-p_{t}^{l})^{2}}$
$\left((p_{t}^{l})^{2}(\eta_{t}^{2}G^{2}+v_{t}^{l})+\frac{U}{K(U-1)}\left(1-\frac{K}{U}\right)4\eta_{t}^{2}G^{2}\right)$
$\leq p_{t}^{l}\left(v_{t}^{l}+\eta_{t}^{2}G^{2}\right)+\frac{1}{(1-p_{t}^{l})^{2}}$
$\left((p_{t}^{l})^{2}(\eta_{t}^{2}G^{2}+v_{t}^{l})+\frac{4\eta_{t}^{2}G^{2}U}{U-1}\right)\sum_{K=1}^{U}P\Big[|\mathcal{U}_{t}^{l}|=K\Big]$
$=p_{t}^{l}\left(v_{t}^{l}+\eta_{t}^{2}G^{2}\right)$
$+\frac{1}{1-p_{t}^{l}}\left((p_{t}^{l})^{2}(\eta_{t}^{2}G^{2}+v_{t}^{l})+\frac{4\eta_{t}^{2}G^{2}U}{U-1}\right)$
$=\frac{p_{t}^{l}}{1-p_{t}^{l}}v_{t}^{l}+\eta_{t}^{2}G^{2}\left(\frac{\frac{4U}{U-1}+p_{t}^{l}}{1-p_{t}^{l}}\right).$
$v_{t+1}^{l}\leq D_{t}^{l}\cdot v_{t}^{l}+E_{t}^{l},\quad v_{1}^{l}=0,$ (A.14)
where
$D_{t}^{l}\triangleq\frac{p_{t}^{l}}{1-p_{t}^{l}}\quad\text{and}\quad E_{t}^{l}\triangleq\eta_{t}^{2}G^{2}\left(\frac{\frac{4U}{U-1}+p_{t}^{l}}{1-p_{t}^{l}}\right).$ (A.15)
Since $T_{t}^{\rm d}\geq T_{t+1}^{\rm d}$, we get $p_{t}^{l}\leq p_{t+1}^{l}$ and $D_{t}^{l}\leq D_{t+1}^{l}$. In addition, since the learning rate satisfies $\eta_{t}\leq 2\eta_{t+1}$, it follows that $\eta_{t}^{2}\leq 4\eta_{t+1}^{2}$ and therefore $E_{t}^{l}\leq 4E_{t+1}^{l}$. Unfolding the recursion yields:
$v_{t+1}^{l}$ $\leq D_{t}^{l}\cdot v_{t}^{l}+E_{t}^{l}\leq D_{t}^{l}D_{t-1}^{l}v_{t-1}^{l}+D_{t}^{l}E_{t-1}^{l}+E_{t}^{l}$
$\leq\sum_{k=1}^{t}E_{k}^{l}\prod_{j=k+1}^{t}D_{j}^{l}+\underbrace{v_{1}^{l}\prod_{k=1}^{t}D_{k}^{l}}_{0}\leq\sum_{k=1}^{t}E_{k}^{l}(D_{t}^{l})^{t-k}$
$\leq\sum_{k=1}^{t}4^{t-k}E_{t}^{l}(D_{t}^{l})^{t-k}=E_{t}^{l}\sum_{k=1}^{t}(4D_{t}^{l})^{t-k}$
$=\frac{1-(4D_{t}^{l})^{t}}{1-4D_{t}^{l}}E_{t}^{l},$
where the last equality holds since, by the conditions of Lemma 3, we have $p_{t}^{l}\leq p_{t}^{1}<0.2$. Substituting this bound into the definition of $D_{t}^{l}$ in Eq. (A.15) implies $D_{t}^{l}<0.25$, and therefore $4D_{t}^{l}<1$. Substituting $D_{t}^{l}$ and $E_{t}^{l}$ given in Eq. (A.15), we have:
$v_{t+1}^{l}$ $=\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big]\leq\frac{1-(4D_{t}^{l})^{t}}{1-4D_{t}^{l}}E_{t}^{l}$
$\leq\frac{E_{t}^{l}}{1-4D_{t}^{l}}=\eta_{t}^{2}G^{2}\left(\frac{\frac{4U}{U-1}+p_{t}^{l}}{1-5p_{t}^{l}}\right).$
Now, summing over the $L$ layers:
$\leq\sum_{l=1}^{L}\eta_{t}^{2}G^{2}\left(\frac{\frac{4U}{U-1}+p_{t}^{l}}{1-5p_{t}^{l}}\right)\leq\eta_{t}^{2}G^{2}\frac{4U}{U-1}\sum_{l=1}^{L}\frac{1+p_{t}^{l}}{1-5p_{t}^{l}}.$
Substituting the bound on $p_{t}^{l}$, as given in Lemma 1, gives:
$\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}\right\|^{2}\right]$
$\leq\eta_{t}^{2}G^{2}\frac{4U}{(U-1)}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}{1-5Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}},$
thus proving Lemma 3. $\square$
A-D Appendix D - Proof of Theorem 1
We start by using the full participation FedAvg update ${\bm{w}}_{t+1}$:
$\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{opt}\|^{2}\Big]$
$=$ $\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}+{\bm{w}}_{t+1}-{\bm{w}}_{opt}\|^{2}\Big]$
$=$ $\underbrace{\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}\|^{2}\Big]}_{S_{1}}+\underbrace{\mathbb{E}\Big[\|{\bm{w}}_{t+1}-{\bm{w}}_{opt}\|^{2}\Big]}_{S_{2}}$
$+$ $\underbrace{2\langle\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1},{\bm{w}}_{t+1}-{\bm{w}}_{opt}\rangle}_{S_{3}}.$
We note that given the batches $\{i_{t}^{u}\}$, $S_{3}$ vanishes by the unbiasedness Lemma 2. $S_{1}$ is bounded using Lemma 3:
$\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}\right\|^{2}\right]$
$\leq$ $\eta_{t}^{2}G^{2}\frac{4U}{(U-1)}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}{1-5Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}.$
The term $S_{2}$ is bounded using Lemma 1 from [18], which states that under Assumption A1 and a step size satisfying $\eta_{t}\leq\frac{1}{4\rho_{s}}$, the following inequality holds:
$\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{opt}\|^{2}\Big]$ $=(1-\eta_{t}\rho_{c})\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{opt}\|^{2}\Big]$ (A.16)
$+\underbrace{\eta_{t}^{2}\mathbb{E}\Big[\|g_{t}-\bar{g}_{t}\|^{2}\Big]}_{\tilde{F}}+6\eta_{t}^{2}\rho_{s}\Gamma,$
where $g_{t}$ and $\bar{g}_{t}$ are defined by:
[…]
$=\eta_{t}^{2}\frac{1}{U^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{S_{t}^{u}}$
$=\eta_{t}^{2}\frac{1}{U^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{\left\lfloor mP_{u}\left(\frac{T_{t}^{\rm d}-B_{u}}{T_{t}^{\rm d}}\right)\right\rfloor}$
$\leq\eta_{t}^{2}\frac{1}{U^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{mP_{u}\left(\frac{T_{t}^{\rm d}-B_{u}}{T_{t}^{\rm d}}\right)-1},$
where the third equality stems from Assumption A2, the forth equality stems from Model Formulation B3 and the inequality from $\frac{1}{\lfloor x\rfloor}\leq\frac{1}{x-1},\quad\forall x>1$.
Combining the bounds for $S_{1}$ and $S_{2}$, we obtain:
$\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{opt}\|^{2}\Big]$ $\leq(1-\eta_{t}\rho_{c})\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{opt}\|^{2}\Big]$
$+\eta_{t}^{2}\left(B_{t}+C_{t}\right),$
where $B_{t}$ and $C_{t}$ are given by:
$B_{t}$ $\triangleq\frac{1}{U^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{mP_{u}\left(\frac{T_{t}^{\rm d}-B_{u}}{T_{t}^{\rm d}}\right)-1}+6\rho_{s}\Gamma$
$C_{t}$ $\triangleq G^{2}\frac{4U}{U-1}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}{1-5Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}.$
$\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{R+1}-{\bm{w}}_{opt}\right\|^{2}\right]$ $\leq\prod_{t=1}^{R}\left(1-\eta_{t}\rho_{c}\right)\Delta_{1}$
$+\sum_{t=1}^{R}\eta_{t}^{2}\left(B_{t}+C_{t}\right)\prod_{\tau=t+1}^{R}\left(1-\eta_{\tau}\rho_{c}\right)$
A-E Appendix E - Proof of Auxiliary Lemma
We derive this expression using integration by parts. For an integer $s>1$:
@@ A-E Appendix E - Proof of Auxiliary Lemma
$Q(s,x)$ $=\left[-\frac{t^{s-1}e^{-t}}{\Gamma(s)}\right]_{t=x}^{\infty}+\frac{1}{\Gamma(s-1)}\int_{x}^{\infty}t^{s-2}e^{-t}dt$
@@
$Q(s,x)=\sum_{k=0}^{s-1}\frac{x^{k}e^{-x}}{k!},$ (A.17)
concluding the proof. $\square$